I have one more from yesterday, I am still stuck, please help.
Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form ax^2+bx+c=0, where a,b,and c are integers(no fractions) solution -3/4,9/4. Thank you.

Answer by sudhanshu_kmr(1152)


x = -3/4 so, 4x+3 = 0
x = 9/4 , 4x-9 = 0
multiply both i.e (4x+3) (4x-9)
you will get.-> 16x^2 -24x -27
so, required equation is : 16x^2 -24x -27 = 0

if any doubt, please contact..